plain wavefront of light of wavelength 8800 A° is incedent on two slites in a screen perpendicular to the direction of light rays if the total separation of 13 fringes on a screen 4m away is 3cm find the distance between slites
1
Expert's answer
2021-01-07T10:55:37-0500

Answer

the distance between slites

"d=\frac{n\lambda D}{x}"

"\frac{13\times8800\times10^{-10} \times4 }{3 \times 10^{-2} }\\=1.5mm"